Simplify a (a^2+ a + 1) + 5 and find its value for (i) a = 0, (ii) a = 1 (iii) a = – 1.

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

a(a^2+a +1)+5
= a* a^2 + a*a + a*1 + 5
=a^3+a^2+a+ 5
(i) putting a= 0 in the equation
we get 0^3+0^2+0+5=5
(ii)putting a=1 in the equation we get
1^3+ 1^2+ 1+5
= 1 + 1 + 1+5 = 8
(iii)Putting a = -1 in the equation we get
(-1)^3+(-1)^2+ (-1)+ 5
= -1 + 1–1+5 = 4

Thanks